### v1

# Look Before You Leap: Problem Elaboration Prompting Improves Mathematical Reasoning in Large Language Models

Abstract
Large language models (LLMs) have exhibited impressive performance across NLP tasks. So far they still face challenges in complex reasoning tasks and can be sensitive to input context. Despite significant efforts have been invested in enhancing reasoning process and improving prefix-prompts robustness, the crucial role of problem context has been overlooked. In this study, we propose a new approach to improve the mathematical capacities of LLMs, named Problem Elaboration Prompting (PEP). Specifically, PEP decomposes and elucidates the problem context before reasoning, thus enhancing the global context modeling and reducing the parsing difficulties. Experiments on datasets demonstrate promising performances on complex reasoning and indicate the beneficial impact for ill-formed problems. For instance, with the GPT-3.5 model (text-davinci-003), we observed a 9.93% improvement with greedy decoding and 8.80% improvement with self-consistency on GSM8k compared to the standard CoT. With ChatGPT (turbo) and PEP, we achieve SOTA performances on SVAMP with 86.2% and GSM8k with 90.98%.
[…]
Look Before You Leap: Problem Elaboration Prompting Improves Mathematical Reasoning in Large Language Models
Haoran Liaothanks: Email: liaohaoran@sjtu.edu.cn , Jidong Tian, Shaohua Hu, Hao He, Yaohui Jin
MoE Key Lab of Artificial Intelligence, AI Institute, Shanghai Jiao Tong University
Abstract content
1.   Introduction
Recent large language models (LLMs), such as the GPT-3 model family with 175 billion parameters (Brown et al., 2020, inter alia), have demonstrated remarkable performance across various NLP tasks. Chain-of-thought (CoT) prompting (Wei et al., 2022b; Kojima et al., 2022) successfully elicits reasoning behavior and emergent abilities (Wei et al., 2022a) by explicitly guiding the model to generate intermediate rationales step by step, further promoting the development of artificial general intelligence (AGI). Despite the success, performing multi-hop and compositional reasoning for complex tasks can still face challenges (Hendrycks et al., 2021; Lewkowycz et al., 2022), even when the required knowledge is limited to the scope of primary school (Cobbe et al., 2021).
Figure 1: Possible inputs and outputs of a LLM, including prefix-prompts, problem context, proposed post-problem elaboration, and step-by-step solutions. The complete results and comparisons are presented in Sec. B. In this work, we adopt a zero-shot setting without using exemplars.
Figure 2: An overview of the proposed Problem Elaboration Prompting (PEP) method with standard prompting and CoT prompting. In specific, PEP combs through the problem context before reasoning, aiming to simplify and clarify the given description and asked question to enhance the language model’s comprehension, thereby facilitating a more clear downstream reasoning process.
One area of research aims to improve the quality of reasoning outputs through diverse rationale decoding strategies (Wang et al., 2022c; Suzgun et al., 2022) and iteration-based answer refinement (Saunders et al., 2022; Kim et al., 2023; Zheng et al., 2023). Considering the sensitivity of the model to inputs (Lu et al., 2022b; Wang et al., 2022a; Shi et al., 2023), another area of research focuses on augmenting the robustness of prefix-prompts (Fu et al., 2022; Wang et al., 2022b; Shao et al., 2023). However, the role of problem has been overlooked. Few works, such as Least-to-Most (L2M) prompting, have paid attention to this part (Zhou et al., 2022a; Khot et al., 2022). But L2M only aims to decompose the final question into sub-questions, ignoring the rest problem description.
Recognizing the known conditions and specified intentions of the problem is imperative for subsequent reasoning. The success of previous efforts that have elicited reasoning capabilities in language models, to some extent, benefits from the implicit parsing of problem context (Perez et al., 2020; Zhou et al., 2022a; Wei et al., 2022b). However, most studies assume that the provided information is concise and relevant, while in real-world situations, the input problem could be ill-formed. The language model can be easily distracted by irrelevant context (Shi et al., 2023) and often struggles to capture intricate implied implicatures (Ruis et al., 2022; Chan et al., 2023). Even when the problem is well-formed, it may still be too complex and unsuitable for the language model’s comprehension. For instance, although the language model can be knowledgeable, it may encounter difficulties in identifying what knowledge are required to answer the question (Bian et al., 2023) or correctly integrating intermediate rationales to generate the overall solution (Press et al., 2022).
In this study, we introduce a new method named Problem Elaboration Prompting (PEP), which involves decomposing and elucidating the problem context prior to reasoning. Our method aims to clarify the problem description and enhance the context modeling. We illustrate the possible inputs and outputs for the LLMs in Fig. 1, showing how problem elaboration improves the reasoning of LLMs. PEP adopts a human-like thought process that emphasizes the importance of thoroughly comprehending the problem’s conditions and requirements before engaging in reasoning: look before you leap. PEP can be easily combined with other prompting methods. In addition, our work is also inspired by previous researches on semantic parsing, which suggest parsing the given problem into specific representations, such as Python code or condensed symbols (Gao et al., 2022; Chen et al., 2022; Hu et al., 2023), to facilitate subsequent reasoning.
We present an overview of PEP in Fig. 2, alongside the standard prompting and CoT prompting. We conduct evaluations of the proposed approach on six reasoning datasets and additionally investigate its performance in addressing the distraction problem. PEP consistently outperforms the standard zero-shot baselines (CoT and PoT (Chen et al., 2022)) across datasets, employing both greedy decoding and self-consistency setting. For instance, with GPT-3.5 (text-davinci-003), we observed a 9.93% improvement with greedy decoding on GSM8k (Cobbe et al., 2021), and 8.80% improvement with self-consistency. When using ChatGPT (Turbo), PEP achieves SOTA with 90.98% on GSM8k, and 86.2% on SVAMP (Patel et al., 2021).
In summary, our contributions are listed below:
  1. 1.
    We propose a new method, Problem Elaboration Prompting (PEP), alongside CoT and self-consistency CoT, for improving LLM reasoning abilities.
  2. 2.
    We evaluate PEP through extensive experimentation using both text-davinci-003 and turbo, exploring the role of problem context. We hope it could inspire future work.
  3. 3.
    We demonstrate the effectiveness of PEP in mitigating the distraction problem, indicating the promising prospect in dealing with ill-formed problems of other types.
2.   Related Works
2.1.   Emergent Abilities and Prompting
With the large-scale unsupervised pre-training technique, LLMs (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023, inter alia) can perform new tasks conditioning on few in-context inputs via mimicking the format of demonstrations (Webson and Pavlick, 2022; Min et al., 2022; Pan et al., 2023). Instruction tuning further improves the generalization of prompts via training a language model to follow the general instructions rather than specific examples (Wei et al., 2021; Ouyang et al., 2022a; Chung et al., 2022).
Instead of directly generating the final answer, chain-of-thought prompting (CoT) methods (Wei et al., 2022b; Creswell et al., 2022; Jung et al., 2022) suggest guiding LLMs to reach the final answer through a step-by-step process, eliciting the emergent reasoning ability (Wei et al., 2022a; Chan et al., 2022; Prystawski and Goodman, 2023). Zelikman et al. (2022) also proposed to use prompts to augment the trainset with rationales. However, such methods typically necessitate an LLM larger than 100B (Wei et al., 2022b), making it difficult to directly apply on small models (Zhang et al., 2023).
2.2.   Improving CoT Reasoning
Various techniques have been proposed to improve the standard CoT (Wang et al., 2022c; Suzgun et al., 2022; Saunders et al., 2022; Gao et al., 2022; Kim et al., 2023; Yao et al., 2023), most with a particular focus on controlling the decoding process and validating intermediate answers. Significant efforts have also been made to improve the prompts selection and ensemble (Fu et al., 2022; Lu et al., 2022a; Zhou et al., 2022b; Shao et al., 2023), to enhance the generation robustness. Besides, increasing the diversity of prompts, questions, or rationales shows notable advantages for the reasoning (Zhang et al., 2022b; Wang et al., 2022b).
Our work is similar to the study conducted by Lampinen et al. (2022), which explores the impact of explanations in context. However, it investigates the post-answer explanation about rationales, while we focus on elaborating the problem context.
2.3.   Collaborating CoT with agents.
Recent works attempt to introduce the concept of agents to enhance the reasoning, explicitly prompting the LLM with external tools (Gao et al., 2022) to play different roles in undertaking decomposed tasks  (Shinn et al., 2023; Park et al., 2023; Wang et al., 2023a). It typically includes planning, cooperation, competition, and interaction with the environment (Xi et al., 2023). Exploring the decision-making process (Jung et al., 2022) with specific agents has also shown promising improvements, including tree-structure (Yao et al., 2023) and graph-structure (McNichols et al., 2023; Cao, 2023).
Our work can also be regarded as a type of LLM-based agent system. Although existing approaches also have the concept of analysis, they focus more on observing the state of the environment and reflecting the feedback (Zheng et al., 2023; Qian et al., 2023; Wu et al., 2023; Miao et al., 2023), while PEP focuses on analyzing and parsing the problem.
3.   Problem Elaboration Prompting
A beneficial cognitive paradigm of humanity is the ability of thoughtful contemplation, as exemplified by a common figurative: look before you leap. In this work, we propose to simulate such a thinking process by prompting the model to elaborate on the problem before generating rationales and answers. In other words, it is crucial for the language model to ensure it has a thorough understanding of the problem space, encompassing the given conditions and the asked question, avoiding a hasty generation and exploration within the answer space (Table 1).
Q: Brian’s friend Bobby has 5 fewer than 3 times as many video games as Brian does. If Brian has 20 video games but lost 5 right before the comparison was made, how many does Bobby have?
Standard CoT:
A: Three times as many video games as Brian has is 3*20=<<3*20=60>>60 video games. Five fewer than 3 times as many video games as Brian has is 60-5=<<60-5=55>>55 video games.
Final answer: 55.
(Wrong: ignore the precondition placed at end.)
PEP + CoT
(+Elaboration; Declaratives):
Segment 1: Bobby has 5 fewer than 3 times as many video games as Brian does. Rephrased: Bobby’s number of video games is 3 times Brian’s number minus 5.
Equation: B = 3B - 5
Segment 2: Brian has 20 video games.
Equation: B = 20
Segment 3: Brian lost 5 video games right before the comparison was made.
Equation: B = 20 - 5
A: Using segment 2’s equation, we know that Brian has 20 video games. Using segment 3’s equation, we know that Brian lost 5 video games, leaving him with 15 video games. We can substitute this value into segment 1’s equation and solve for Bobby’s number of video games. B = 3(15) - 5, B = 40.
Final answer: 40.
(Correct: process conditions in suitable order.)
Table 1: An example of PEP with turbo. After capturing the relevant conditions presented upfront, the vanilla CoT wrongly engages in hasty reasoning (red), neglecting the important prerequisite conditions placed afterwards (cyan). Meanwhile, PEP benefits from the preprocessing elaboration and composites conditions in the correct order (blue).
3.1.   Method Formulation
Given the problem context denoted by $P=(X,y_{0})\in D$, where $X=\{x_{1},x_{2},\cdots,x_{i}\}$ represent the statements and $y_{0}$ represents the asked question, prompting methods attempt to leverage the strength of LLMs to solve the task by generating rationale $y_{i}=f_{M}(X,y_{0};p_{t},\hat{P}_{k}|y_{<i})$ step by step until reaching the final answer $y_{t}$. Specifically, $p_{t}$ denotes zero-shot instructions and $\hat{P}_{k}$ indicate $k$ concatenated exemplars $(\hat{X}_{k},\hat{y}_{k,0})$. Note that such prompting methods do not modify the LLM $M$.
In the PEP-aid language model, we suggest to pre-process the given problem $P$ by decomposing and elucidating the context into smaller and more concise segments to enhance LLMs’ comprehension: $P^{\prime}=(\{x_{1}^{\prime},x_{2}^{\prime},\cdots,x_{m}^{\prime}\},y_{0}^{\prime})=f_{M}(X,y_{0};p_{e})$, where $m\geq i$ and $p_{e}$ is a specific instruction. Then, the LLM can continue its reasoning by $y_{i}=f_{M}(X,y_{0};p_{t},\hat{P}_{k}|P^{\prime},y_{<i})$ utill reaching $y_{t}$. Thus, PEP can easily be combined with previous prompting methods.
3.2.   Designing Principles
The design principle of problem elaboration in this study consists of two aspects: (i) decomposing: breaking down the original sentences into distinct segments to disentangle complex or intertwined concepts; (ii) elucidating: providing explanations or rephrasing the segments in a manner that is more conducive to the model’s understanding.
The concept of decomposition is widely spread in many previous works. For example, the L2M prompting (Zhou et al., 2022a) suggests breaking down the asked questions into sub-questions and then prompts the model to solve them one by one. The Plan-and-Solve prompting (Wang et al., 2023b) suggests generating an overall plan before reasoning. In contrast, PEP takes a different approach by breaking down the entire problem into simpler segments, rather than just decomposing the final question. Furthermore, PEP does not require deducing a plan or inferring the sub-questions. Instead, it focuses on organizing and clarifying the existing information from the problem.
Meanwhile, it could be beneficial to elucidate the segments following decomposition, as it elicits the model to organize existing information in a comprehensive view of the problem and recognize the underlying implicatures of the question. Moreover, it introduces diversity into the context, which has been demonstrated to enhance reasoning (Zhang et al., 2022b; Wang et al., 2022b), thereby mitigating the risk of relying on specific words or descriptions as shortcuts. A break-down analysis of our designed principles is presented in Sec. 4.4.
Q/A pair Declaratives Interrogative Mixed
19.5 47.0 23.0 10.5
Table 2: Distribution of PEP forms. We annotated the corresponding types in respective cases.
3.3.   Instruction Generation
We utilize the zero-shot setting to prompt the language models since the elaboration of different problems can be diverse. We apply LLM to generate alternatives. Then, the instructions are evaluated on 200 points randomly sampled from the GSM8k trainset. The final selected instruction is “Decompose the given question into smaller segments, elucidating each segment as you rephrase it.” The details of the selection process and all used prompts are provided in the Appendix.
To investigate the behavior of PEP, we randomly selected 200 instances from the experiments with a manual categorization. PEP primarily utilizes question-answer pairs, declarative sentences, and interrogative sentences to review and examine the problem context (see Table 2). There are around 10.5% of instances that can be mixed with sub-questions, planning instructions, or intertwined rationales. We find it primarily occurs when the questions themselves contain instructions or options.
4.   Experiment
4.1.   Setup
Dataset #Test #AT domain
SingleEq 508 free math
GSM8k 1319 free math
AQuA 254 options math
SVAMP 1000 free math
MMLU-M 500 options math
GSMIC-1k 1000 free distraction
Table 3: Statistical information of used datasets. #Test refers to the number of test samples and #AT denotes the answer types.
Datasets.
We evaluate PEP on four elementary datasets, with a focus on mathematical reasoning: (1) SingleEq (Koncel-Kedziorski et al., 2015), (2) GSM8k (Cobbe et al., 2021), (3) SVAMP (Patel et al., 2021), (4) AQuA (Ling et al., 2017); and five challenging subsets from MMLU-STEM Hendrycks et al. (2020), denoted by (5) MMLU-MATH. Furthermore, we investigate the distraction problem using GSMIC (Shi et al., 2023): we randomly sampled 500 examples separately for 2-step problems and m-step problems, denoted by (6) GSMIC-1k. Statistical information are presented in Table 3.
Dataset
Model Prompt PEP (ours) SingleEq GSM8k AQuA SVAMP Average
CoT (Kojima et al., 2022) 87.40 50.19 38.19 72.1 61.97
90.35 60.12 40.55 75.7 66.68
(+2.95) (+9.93) (+2.36) (+3.6) (+4.71)
PoT (Chen et al., 2022) 95.67 63.84 29.13 78.3 66.74
96.46 64.75 29.53 81.3 68.01
GPT-3.5 text-davinci-003 (+0.79) (+0.91) (+0.40) (+3.0) (+1.27)
CoT (Kojima et al., 2022) 91.93 80.44 55.12 78.8 76.57
91.93 78.92 57.09 79.8 76.94
(+0.0) (-1.52) (+1.97) (+1.0) (+0.37)
PoT (Chen et al., 2022) 97.05 75.82 36.61 81.6 72.77
96.06 76.57 37.80 84.1 73.63
ChatGPT gpt-3.5- turbo-0301 (-0.99) (+0.75) (+1.19) (+2.5) (+0.86)
Table 4: Accuracies (x100) of our approach (PEP) against zero-shot CoT and zero-shot PoT. All results are with greedy decoding (GD), i.e., the temperature is set to zero. PEP demonstrate an overall improvements compared to baselines, especially with davinci on GSM8k ($50.19\%\to 60.12\%$).
Dataset
Model Prompt PEP (ours) SingleEq GSM8k AQuA SVAMP Average
CoT+SC (Wang et al., 2022c) 87.99 52.31 39.76 71.7 62.94
89.57 61.11 39.76 73.6 66.01
GPT-3.5 text-davinci-003 (+1.58) (+8.80) (+0.0) (1.90) (+3.07)
CoT+SC (Wang et al., 2022c) 93.50 88.70 69.29 84.6 84.02
93.50 90.98 68.11 86.2 84.70
ChatGPT gpt-3.5- turbo-0301 (+0.0) (+2.28) (-1.18) (+1.6) (+0.68)
Table 5: Solve rates (x100) of our approach (PEP) against standard prompting (zero-shot CoT). All results are with self-consistency (SC), where temperature T=0.7 and reasoning paths are sampled for 20 times.
Baselines & Prompts.
We compare PEP against (1) zero-shot CoT (Kojima et al., 2022) for textual reasoning, and (2) zero-shot PoT (Chen et al., 2022) for code-based reasoning. We also compare (3) Irr-Inst. that Shi et al. (2023) suggests to mitigate distraction problem. We added some conjunctions and adjusted the template of PoT to ensure the format is valid. All prompts are presented in the Appendix in detail.
Language Models & Decoding
We conduct our main experiments on (1) the “text-davinci-003” model  (denoted as davinci hereafter) and (2) the “gpt-3.5-turbo-0301” model (denoted as turbo hereafter) (Brown et al., 2020; Ouyang et al., 2022b; OpenAI, 2023). Both models are employed via OpenAI API. We evaluate the performance with both greedy decoding and self-consistency decoding (SC) (Wang et al., 2022c) to ensure the reproducibility of experiments. There may be variations in the results obtained from other versions or updated models.
4.2.   Main Results
The main results are presented in Table 4 and Table 5. Considering the token cost, we only compared PEP with CoT using self-consistency. Overall, our proposed PEP performs better compared to the vanilla CoT and PoT, achieving SOTAs with turbo (90.98% on GSM8k and 86.2% on SVAMP). We also evaluate PEP against vanilla CoT on challenging MMLU-MATH, which includes five more complex mathematical tasks, as shown in Fig. 3. PEP presents consistently improvements for most subtopics with both davinci ($\xrightarrow{}$44.2%) and turbo ($\xrightarrow{}$56.0%).
Figure 3: Performances on MMLU-Math. All results are generated with greedy search. Although there is a decline for the HighSchoolMath subtopic, PEP exhibits an overall enhancement compared to the vanilla CoT across the complex subtopics.
Micro Accuracy Macro Accuracy
Method (turbo-0301) 2 steps >2 steps Overall Norm 2 steps >2 steps Overall Norm
CoT 83.2 74.2 78.7 85.42 83.26 72.05 77.65 84.41
PEP 83.6 75.6 79.6 86.4 83.11 73.99 78.55 85.38
(+0.4) (+1.4) (+0.9) (+0.98) (-0.15) (+1.95) (+0.90) (+0.98)
+ Irr-Inst. (Shi et al., 2023)
CoT 85.6 80.8 83.2 90.3 86.28 79.19 82.74 89.93
PEP 87.4 84.0 85.7 93.02 86.63 83.83 85.23 92.65
(+1.8) (+3.2) (+2.5) (+2.71) (+0.35) (+4.65) (+2.5) (+2.72)
Table 6: Micro and macro accuracies (×100) on GSMIC-1k. We follow the metrics and instruction suggested in (Shi et al., 2023). Norm is the overall accuracy normalized by the scores of base problems, measuring how a method is affected by the distractors. All results are generated with greedy decoding.
Specifically, when working with davinci, which may be less capable in problem recognition compared to turbo, PEP shows significant improvement and outperforms the baselines by a substantial margin (9.93% and 8.80% on GSM8k for two decoding settings separately). Under greedy decoding, PEP exhibits an overall superiority to both CoT and PoT. The enhancement for davinci is more pronounced with CoT, whereas for turbo, it is PoT. This could suggest that PEP does not have a particular reliance on the manner of downstream reasoning. In addition, PEP with turbo shows a decrease on GSM8k under GD, but performs better with SC. In Table 5, PEP achieves the same accuracy as CoT for some cases and performs worse on AQuA. To further analyze the performance of PEP with SC, we provide a break-down analysis in the Sec. 4.4.
4.3.   Distraction Problem
LLMs may encounter difficulties when handling ill-formed problems. One particular challenge is known as the distraction problem (Shi et al., 2023), which means when adding irrelevant sentences to a given problem, the model can be easily distracted and further generate erroneous results. These additional sentences can be completely irrelevant or relevant to the problem but have no impact on the reasoning. An example is in Table 7, comparing the performances of CoT and PEP when dealing with the distracting problem.
In this study, we investigated PEP on distraction problem based on GSMIC dataset (Shi et al., 2023), which is constructed from a base subset of GSM8k (Cobbe et al., 2021). The irrelevant sentences are randomly added according to certain templates and pre-designed distracting factors. Two metrics are used in our experiments: (1) micro-accuracy: averaged accuracy per example, and (2) macro accuracy: averaged accuracy per base problem. The main results are shown in Table 6.
Figure 4: Break-down analysis of the number of self-consistent sampling paths. Comparing with the results in Table 5, we can find PEP presents consistently better performances than the standard CoT for most cases.
Q: The average number of fruits per basket in five baskets is 25. Ada bought 20 tomatoes from the grocery store. If basket A has 15 apples, B has 30 mangoes, C has 20 peaches, D has 25 pears and E has some bananas, how many bananas are in basket E?
Standard CoT (turbo with greedy decoding):
A: 1. Total number of fruits: 25 x 5 = 125. 2. Add the tomatoes to the total number of fruits: 125 + 20 = 145. 3. Subtract the number of fruits in baskets A, B, C, and D from the total number of fruits: Bananas in basket E = 145 - 90 = 55.  (Wrong: distracted.)
PEP + CoT (turbo with greedy decoding):
(+Elaboration; Interrogative):
Segment 1: What is the average number of fruits per basket in five baskets?
Segment 2: How many tomatoes did Ada buy from the grocery store?
Segment 3: What is the fruit distribution in each basket?
Segment 4: How many bananas are in basket E?
A: Segment 1: The average number is 25.
Segment 2: Ada bought 20 tomatoes.
Segment 3: A: 15 apples, B: 30 mangoes, C: 20 peaches, D: 25 pears, E: some bananas.
Segment 4: To find out how many bananas are in basket E, we need to use the information given in Segment 1 and Segment 3.
Total number is 25 x 5 = 125. In A, B, C, and D is 15 + 30 + 20 + 25 = 90. Therefore, the number of fruits in basket E is 125 - 90 = 35.  (Correct: concentrated.)
Table 7: When adding irrelevant sentence (cyan) to the problem, CoT is easily distracted (red), while PEP is more focused (blue). We omitted and rephrased certain sentences for clarity.
PEP can help alleviate the distraction problem.
In Table 6, PEP outperforms CoT for both metrics, demonstrating better robustness to distraction problem. We can further observe that PEP performs better on m-step problems than 2-step tasks, indicating a more significant improvement where the premises are intricately coupled with irrelevant factors. The macro accuracy of PEP declines for 2-step problems, indicating the performance of PEP may vary for different base problems.
PEP performs better when prompted with prior knowledge.
When the model is consciously prompted to ignore irrelevant content for the given problem, referred to as Irr-Inst. (Shi et al., 2023), we observed significant improvements in both CoT and PEP. However, PEP exhibits more substantial and consistent enhancements compared to the CoT, with a larger margin ($0.98\%\to 2.71\%$). It is promising that PEP may also be effective for other types of ill-formed problems with prior knowledge.
4.4.   Ablation Study
Method AQuA GSM8k GSMIC-1k
CoT 55.12 80.44 78.7
PEP 57.09 78.92 79.6
- Dec 53.15 78.17 76.7
- Elu 62.6 78.70 78.6
EtD 54.33 76.27 77.1
Table 8: Ablation Study. Dec and Elu indicate prompting the model to only perform decomposition and perform elucidation, respectively. EtD means prompting the model to elucidate then decompose.
In this section, we conducted ablation studies to verify the criticality of PEP principles, performances of different SC sampling settings, and the effect of diverse distracting factors.
Micro Accuracy Macro Accuracy
Topic Role Overlap Num. Range Topic Role Overlap Num. Range
Method (turbo) In Off Yes No In Out In Off Yes No In Out
CoT 68.58 87.71 74.95 81.94 78.43 78.97 66.86 87.15 76.49 81.72 78.68 78.44
PEP 73.67 84.88 79.27 79.89 78.43 80.75 69.9 86.11 78.23 80.45 78.01 80.04
(+5.1) (-2.84) (+4.32) (-2.05) (+0.0) (+1.79) (+3.04) (-1.04) (+1.73) (-1.27) (-0.67) (+1.6)
+ Irr-Inst. (Shi et al., 2023)
CoT 75.58 89.98 83.8 82.68 83.27 83.13 74.75 89.84 84.46 82.24 82.48 83.4
PEP 79.19 91.49 85.53 85.85 83.06 88.29 77.47 91.76 84.61 86.16 82.61 86.98
(+3.61) (+1.51) (+1.73) (+3.17) (-0.2) (+5.16) (+2.72) (+1.92) (+0.14) (+3.92) (+0.13) (+3.58)
Table 9: The breakdown accuracies (×100) w.r.t. the added irrelevant sentence factors. Lower accuracy suggests the model is more sensitive to irrelevant contexts of that type.
Example 1 (PEP misleads implicit meaning)
Example 2 (PEP discards important information)
Q: An interior design firm offers installation for \$129.00. It includes hanging 4 mirrors, 2 shelves, 1 chandelier, and 10 pictures. They will install additional items for an extra \$15.00 per item. Angela has 6 mirrors and 2 chandeliers and 20 pictures that she needs installed/hung. How much will this cost her?
Q: Mrs. Harrington bought 12 boxes of oranges at the market. She gave her mom and her sister 2 boxes of oranges each. Then she kept 1/4 of the oranges and sold the rest. How many oranges did Mrs. Harrington sell if each box contains 20 oranges?
PEP + CoT; Mixed:
PEP + CoT; Q/A pairs:
S1: What is the base installation fee for the interior design firm? \$129.00. S2: What is included in the base installation fee? 4 mirrors, 2 shelves, 1 chandelier, and 10 pictures. S3: How much does the interior design firm charge for each additional item installed? Extra \$15.00 per item installed. S4: How many additional items does Angela need installed? Angela needs 6 mirrors, 2 chandeliers, and 20 pictures installed, which is a total of 28 additional items.
Solution: The cost of the additional items is 28 items x \$15.00 per item = \$420.00. Therefore, the total cost for Angela’s installation is \$129.00 + \$420.00 = \$549.00.
S1: How many boxes of oranges did Mrs. Harrington buy at the market? 12 boxes. S2: How many boxes of oranges did Mrs. Harrington give to her mom and her sister? 2 boxes. S3: How many boxes of oranges did Mrs. Harrington keep for herself? Mrs. Harrington kept 1/4 of the oranges.S4: How many boxes of oranges did Mrs. Harrington sell? The rest.
Solution: 1. Calculate the total number of oranges Mrs. Harrington bought: 12x20=240. 2. Calculate the total number of oranges she gave to her mom and her sister: 2x2=4, 4x20=80. 3. Total number of oranges kept for herself: 1/4x240=60. 4. Calculate the total number of oranges sold: 240-80-60=100.
Table 10: Examples for Error Analysis using turbo (GD). When pre-processing the problems, PEP may changes the underline meaning and discarding information (red) of the original sentence (cyan).
Break-down analysis of PEP components.
We verify two components proposed in PEP, namely (1) Dec: decomposing only and (2) Elu: elucidating only. The results are presented in Table 8. It can be observed that both Dec and Elu are crucial for PEP. Besides, Elu outperforms both CoT and PEP on AQuA, indicating that the effectiveness of PEP’s components may vary depending on the context. For instance, the problem may be challenging but have already been composed of simple segments. Additionally, we evaluated the prompt of elucidating the problem before decomposition (denoted as EtD). EtD consistently performs worse than PEP and CoT. This may be because elucidation first increases the difficulty of decomposition.
Break-down analysis of different sampled reason paths.
To further investigate PEP and CoT with self-consistent voting, we conducted a break-down analysis with different numbers of reasoning paths (as shown in Fig. 4). Our results show that, for both davinci and turbo, PEP outperforms CoT in most cases. However, increasing the path number does not always lead to improved performances.
Breakdown analysis of distracting factors.
We analyzed the performance of CoT and PEP using different distracting factors, shown in Table 9. It can be observed that our proposed PEP performs better in handling of (1) in-topic sentence with, (2) role name overlap, and (3) out-range numbers, indicating the effectiveness of PEP in recognizing and parsing the problem context. PEP may be confused by the completely irrelevant context. When prompted with prior knowledge (Irr-Inst.) to ignore irrelevant sentences in the input context, PEP consistently outperforms standard CoT by a larger margin on most distracting factors.
4.5.   Error Analysis
To further analyze the behavior of PEP, we conduct error analysis in Table 10. Although PEP only focuses on the problem, prompting the model to perform easier decomposition and elucidation , errors may still occur. When elaborating the problem, PEP may ignore the potential implicatures of the original sentence, resulting in ambiguity of rephrasing. In addition, PEP may focus too much on the local clause and overlook the nested logical structure and temporal relations for the given conditions.
Except increasing the cost of context length, PEP may be inefficient for very long problem descriptions. For certain forms of data, such as short but challenging questions, structured data in table, it could be difficult to elaborate. PEP may break the continuous context, thus changing the implicit implicatures. Besides, PEP may focus too much on the locality, constraining the associative thinking. We present these error cases in the Appendix.
5.   Conclusion
We propose a new method, named Problem Elaboration Prompting (PEP), for improving the reasoning capacities of LLMs. PEP offers several advantages: 1) it improves reasoning compared to CoT and achieved SOTAs for several datasets; 2) it help alleviate the distraction problem, indicating the promise in dealing with ill-formed problems of other types; 3) it requires no extra inference for plans or sub-questions. Thus, it can be combined with other methods that focus on prefix-prompts robustness and rationale qualities.
To further enhance PEP, future research could explore combining more prior knowledge, designing better prompts, and standardizing LLM’s behaviors. We hope this could inspire future exploration of the role of problem context.
Limitations
Method limitations.
As discussed in Sec. 4.5, PEP could introduce errors and biases into the context and faces difficulty to apply on small models (Touvron et al., 2023) as CoT (Zhang et al., 2023). We present more error cases in the Appendix. Moreover, the generation and selection strategies of prompts are vanilla. Investing more resources into searching, like choosing instruction for each dataset individually, may further improve the metrics.
Method coverage.
As PEP is the first prompting method known to us that pre-processes the problem context before reasoning, we only compared it with the zero-shot CoT, PoT, SC-CoT and Irr-Inst.. Our method can potentially be combined with other prompting methods that focus on non-problematic parts for future work.
Risks and ethics.
While we have not identified any immediate risks, the use of PEP may pose ethical concerns that require careful consideration in the future.
6.   Bibliographical References
[…]
  • Bian et al. (2023) Ning Bian, Xianpei Han, Le Sun, Hongyu Lin, Yaojie Lu, and Ben He. 2023. Chatgpt is a knowledgeable but inexperienced solver: An investigation of commonsense problem in large language models. ArXiv, abs/2303.16421.
  • Brown et al. (2020) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al. 2020. Language models are few-shot learners. Advances in neural information processing systems, 33:1877–1901.
  • Cao (2023) Lang Cao. 2023. Enhancing reasoning capabilities of large language models: A graph-based verification approach. ArXiv.
  • Chan et al. (2023) Chunkit Chan, Jiayang Cheng, Weiqi Wang, Yuxin Jiang, Tianqing Fang, Xin Liu, and Yangqiu Song. 2023. Chatgpt evaluation on sentence level relations: A focus on temporal, causal, and discourse relations. ArXiv, abs/2304.14827.
  • Chan et al. (2022) Stephanie Chan, Adam Santoro, Andrew Lampinen, Jane Wang, Aaditya Singh, Pierre Richemond, James McClelland, and Felix Hill. 2022. Data distributional properties drive emergent in-context learning in transformers. Advances in Neural Information Processing Systems, 35:18878–18891.
  • Chen et al. (2022) Wenhu Chen, Xueguang Ma, Xinyi Wang, and William W. Cohen. 2022. Program of thoughts prompting: Disentangling computation from reasoning for numerical reasoning tasks. ArXiv, abs/2211.12588.
  • Chowdhery et al. (2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, Parker Schuh, Kensen Shi, Sasha Tsvyashchenko, Joshua Maynez, Abhishek Rao, Parker Barnes, Yi Tay, Noam M. Shazeer, Vinodkumar Prabhakaran, Emily Reif, Nan Du, Benton C. Hutchinson, Reiner Pope, James Bradbury, Jacob Austin, Michael Isard, Guy Gur-Ari, Pengcheng Yin, Toju Duke, Anselm Levskaya, Sanjay Ghemawat, Sunipa Dev, Henryk Michalewski, Xavier García, Vedant Misra, Kevin Robinson, Liam Fedus, Denny Zhou, Daphne Ippolito, David Luan, Hyeontaek Lim, Barret Zoph, Alexander Spiridonov, Ryan Sepassi, David Dohan, Shivani Agrawal, Mark Omernick, Andrew M. Dai, Thanumalayan Sankaranarayana Pillai, Marie Pellat, Aitor Lewkowycz, Erica Moreira, Rewon Child, Oleksandr Polozov, Katherine Lee, Zongwei Zhou, Xuezhi Wang, Brennan Saeta, Mark Díaz, Orhan Firat, Michele Catasta, Jason Wei, Kathleen S. Meier-Hellstern, Douglas Eck, Jeff Dean, Slav Petrov, and Noah Fiedel. 2022. Palm: Scaling language modeling with pathways. ArXiv, abs/2204.02311.
  • Chung et al. (2022) Hyung Won Chung, Le Hou, S. Longpre, Barret Zoph, Yi Tay, William Fedus, Eric Li, Xuezhi Wang, Mostafa Dehghani, Siddhartha Brahma, Albert Webson, Shixiang Shane Gu, Zhuyun Dai, Mirac Suzgun, Xinyun Chen, Aakanksha Chowdhery, Dasha Valter, Sharan Narang, Gaurav Mishra, Adams Wei Yu, Vincent Zhao, Yanping Huang, Andrew M. Dai, Hongkun Yu, Slav Petrov, Ed Huai hsin Chi, Jeff Dean, Jacob Devlin, Adam Roberts, Denny Zhou, Quoc V. Le, and Jason Wei. 2022. Scaling instruction-finetuned language models. ArXiv, abs/2210.11416.
  • Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Jacob Hilton, Reiichiro Nakano, Christopher Hesse, and John Schulman. 2021. Training verifiers to solve math word problems. ArXiv, abs/2110.14168.
  • Creswell et al. (2022) Antonia Creswell, Murray Shanahan, and Irina Higgins. 2022. Selection-inference: Exploiting large language models for interpretable logical reasoning. ArXiv, abs/2205.09712.
  • Fu et al. (2022) Yao Fu, Hao-Chun Peng, Ashish Sabharwal, Peter Clark, and Tushar Khot. 2022. Complexity-based prompting for multi-step reasoning. ArXiv, abs/2210.00720.
  • Gao et al. (2022) Luyu Gao, Aman Madaan, Shuyan Zhou, Uri Alon, Pengfei Liu, Yiming Yang, Jamie Callan, and Graham Neubig. 2022. Pal: Program-aided language models. ArXiv, abs/2211.10435.
  • Hendrycks et al. (2020) Dan Hendrycks, Collin Burns, Steven Basart, Andy Zou, Mantas Mazeika, Dawn Song, and Jacob Steinhardt. 2020. Measuring massive multitask language understanding. In International Conference on Learning Representations.
  • Hendrycks et al. (2021) Dan Hendrycks, Collin Burns, Saurav Kadavath, Akul Arora, Steven Basart, Eric Tang, Dawn Xiaodong Song, and Jacob Steinhardt. 2021. Measuring mathematical problem solving with the math dataset. ArXiv, abs/2103.03874.
  • Hu et al. (2023) Hanxu Hu, Hongyuan Lu, Huajian Zhang, Wai Lam, and Yue Zhang. 2023. Chain-of-symbol prompting elicits planning in large langauge models. ArXiv, abs/2305.10276.
  • Jung et al. (2022) Jaehun Jung, Lianhui Qin, Sean Welleck, Faeze Brahman, Chandra Bhagavatula, Ronan Le Bras, and Yejin Choi. 2022. Maieutic prompting: Logically consistent reasoning with recursive explanations. In Conference on Empirical Methods in Natural Language Processing.
  • Khot et al. (2022) Tushar Khot, H. Trivedi, Matthew Finlayson, Yao Fu, Kyle Richardson, Peter Clark, and Ashish Sabharwal. 2022. Decomposed prompting: A modular approach for solving complex tasks. ArXiv, abs/2210.02406.
  • Kim et al. (2023) Geunwoo Kim, Pierre Baldi, and Stephen McAleer. 2023. Language models can solve computer tasks. ArXiv, abs/2303.17491.
  • Kojima et al. (2022) Takeshi Kojima, Shixiang Shane Gu, Machel Reid, Yutaka Matsuo, and Yusuke Iwasawa. 2022. Large language models are zero-shot reasoners. Advances in Neural Information Processing Systems.
  • Koncel-Kedziorski et al. (2015) Rik Koncel-Kedziorski, Hannaneh Hajishirzi, Ashish Sabharwal, Oren Etzioni, and Siena Dumas Ang. 2015. Parsing algebraic word problems into equations. Transactions of the Association for Computational Linguistics, 3:585–597.
  • Koralus and Wang-Ma’scianica (2023) Philipp E. Koralus and Vincent Wang-Ma’scianica. 2023. Humans in humans out: On gpt converging toward common sense in both success and failure. ArXiv, abs/2303.17276.
  • Lampinen et al. (2022) Andrew Kyle Lampinen, Ishita Dasgupta, Stephanie C. Y. Chan, Kory Matthewson, Michael Henry Tessler, Antonia Creswell, James L. McClelland, Jane X. Wang, and Felix Hill. 2022. Can language models learn from explanations in context? In Conference on Empirical Methods in Natural Language Processing.
  • Lewkowycz et al. (2022) Aitor Lewkowycz, Anders Johan Andreassen, David Dohan, Ethan Dyer, Henryk Michalewski, Vinay Venkatesh Ramasesh, Ambrose Slone, Cem Anil, Imanol Schlag, Theo Gutman-Solo, et al. 2022. Solving quantitative reasoning problems with language models. In Advances in Neural Information Processing Systems.
  • Ling et al. (2017) Wang Ling, Dani Yogatama, Chris Dyer, and Phil Blunsom. 2017. Program induction by rationale generation: Learning to solve and explain algebraic word problems. In Annual Meeting of the Association for Computational Linguistics.
  • Lu et al. (2022a) Pan Lu, Liang Qiu, Kai-Wei Chang, Ying Nian Wu, Song-Chun Zhu, Tanmay Rajpurohit, Peter Clark, and A. Kalyan. 2022a. Dynamic prompt learning via policy gradient for semi-structured mathematical reasoning. ArXiv, abs/2209.14610.
  • Lu et al. (2022b) Yao Lu, Max Bartolo, Alastair Moore, Sebastian Riedel, and Pontus Stenetorp. 2022b. Fantastically ordered prompts and where to find them: Overcoming few-shot prompt order sensitivity. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 8086–8098.
  • McNichols et al. (2023) Hunter McNichols, Wanyong Feng, Jaewook Lee, Alexander Scarlatos, Digory Smith, Simon Woodhead, and Andrew Lan. 2023. Exploring automated distractor and feedback generation for math multiple-choice questions via in-context learning. ArXiv.
  • Miao et al. (2023) Ning Miao, Yee Whye Teh, and Tom Rainforth. 2023. Selfcheck: Using llms to zero-shot check their own step-by-step reasoning. ArXiv.
  • Min et al. (2022) Sewon Min, Xinxi Lyu, Ari Holtzman, Mikel Artetxe, Mike Lewis, Hannaneh Hajishirzi, and Luke Zettlemoyer. 2022. Rethinking the role of demonstrations: What makes in-context learning work? In Conference on Empirical Methods in Natural Language Processing.
  • OpenAI (2023) OpenAI. 2023. Gpt-4 technical report. ArXiv, abs/2303.08774.
  • Ouyang et al. (2022a) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022a. Training language models to follow instructions with human feedback. Advances in Neural Information Processing Systems, 35:27730–27744.
  • Ouyang et al. (2022b) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022b. Training language models to follow instructions with human feedback. Advances in Neural Information Processing Systems, 35:27730–27744.
  • Pan et al. (2023) Jane Pan, Tianyu Gao, Howard Chen, and Danqi Chen. 2023. What in-context learning "learns" in-context: Disentangling task recognition and task learning. ArXiv, abs/2305.09731.
  • Park et al. (2023) Joon Sung Park, Joseph C. O’Brien, Carrie J. Cai, Meredith Ringel Morris, Percy Liang, and Michael S. Bernstein. 2023. Generative agents: Interactive simulacra of human behavior. ArXiv.
  • Patel et al. (2021) Arkil Patel, S. Bhattamishra, and Navin Goyal. 2021. Are nlp models really able to solve simple math word problems? In North American Chapter of the Association for Computational Linguistics.
  • Perez et al. (2020) Ethan Perez, Patrick Lewis, Wen tau Yih, Kyunghyun Cho, and Douwe Kiela. 2020. Unsupervised question decomposition for question answering. In Conference on Empirical Methods in Natural Language Processing.
  • Press et al. (2022) Ofir Press, Muru Zhang, Sewon Min, Ludwig Schmidt, Noah A. Smith, and Mike Lewis. 2022. Measuring and narrowing the compositionality gap in language models. ArXiv, abs/2210.03350.
  • Prystawski and Goodman (2023) Ben Prystawski and Noah D. Goodman. 2023. Why think step-by-step? reasoning emerges from the locality of experience. ArXiv, abs/2304.03843.
  • Qian et al. (2023) Cheng Qian, Chi Han, Yi R Fung, Yujia Qin, Zhiyuan Liu, and Heng Ji. 2023. Creator: Disentangling abstract and concrete reasonings of large language models through tool creation. ArXiv.
  • Ruis et al. (2022) Laura Ruis, Akbir Khan, Stella Rose Biderman, Sara Hooker, Tim Rocktaschel, and Edward Grefenstette. 2022. Large language models are not zero-shot communicators. ArXiv, abs/2210.14986.
  • Saunders et al. (2022) William Saunders, Catherine Yeh, Jeff Wu, Steven Bills, Long Ouyang, Jonathan Ward, and Jan Leike. 2022. Self-critiquing models for assisting human evaluators. ArXiv, abs/2206.05802.
  • Shao et al. (2023) Zhihong Shao, Yeyun Gong, Yelong Shen, Minlie Huang, Nan Duan, and Weizhu Chen. 2023. Synthetic prompting: Generating chain-of-thought demonstrations for large language models. ArXiv, abs/2302.00618.
  • Shi et al. (2023) Freda Shi, Xinyun Chen, Kanishka Misra, Nathan Scales, David Dohan, Ed Huai hsin Chi, Nathanael Scharli, and Denny Zhou. 2023. Large language models can be easily distracted by irrelevant context. ArXiv, abs/2302.00093.
  • Shinn et al. (2023) Noah Shinn, Federico Cassano, Beck Labash, Ashwin Gopinath, Karthik Narasimhan, and Shunyu Yao. 2023. Reflexion: Language agents with verbal reinforcement learning.
  • Suzgun et al. (2022) Mirac Suzgun, Luke Melas-Kyriazi, and Dan Jurafsky. 2022. Follow the wisdom of the crowd: Effective text generation via minimum bayes risk decoding. ArXiv, abs/2211.07634.
  • Touvron et al. (2023) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, et al. 2023. Llama: Open and efficient foundation language models. arXiv preprint arXiv:2302.13971.
  • Wang et al. (2022a) Boshi Wang, Sewon Min, Xiang Deng, Jiaming Shen, You Wu, Luke Zettlemoyer, and Huan Sun. 2022a. Towards understanding chain-of-thought prompting: An empirical study of what matters. ArXiv, abs/2212.10001.
  • Wang et al. (2023a) Guanzhi Wang, Yuqi Xie, Yunfan Jiang, Ajay Mandlekar, Chaowei Xiao, Yuke Zhu, Linxi Fan, and Anima Anandkumar. 2023a. Voyager: An open-ended embodied agent with large language models. ArXiv.
  • Wang et al. (2023b) Lei Wang, Wanyu Xu, Yihuai Lan, Zhiqiang Hu, Yunshi Lan, Roy Ka-Wei Lee, and Ee-Peng Lim. 2023b. Plan-and-solve prompting: Improving zero-shot chain-of-thought reasoning by large language models. ArXiv, abs/2305.04091.
  • Wang et al. (2022b) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Huai hsin Chi, and Denny Zhou. 2022b. Rationale-augmented ensembles in language models. ArXiv, abs/2207.00747.
  • Wang et al. (2022c) Xuezhi Wang, Jason Wei, Dale Schuurmans, Quoc Le, Ed Huai hsin Chi, and Denny Zhou. 2022c. Self-consistency improves chain of thought reasoning in language models. ArXiv, abs/2203.11171.
  • Webson and Pavlick (2022) Albert Webson and Ellie Pavlick. 2022. Do prompt-based models really understand the meaning of their prompts? In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 2300–2344.
  • Wei et al. (2021) Jason Wei, Maarten Bosma, Vincent Zhao, Kelvin Guu, Adams Wei Yu, Brian Lester, Nan Du, Andrew M Dai, and Quoc V Le. 2021. Finetuned language models are zero-shot learners. In International Conference on Learning Representations.
  • Wei et al. (2022a) Jason Wei, Yi Tay, Rishi Bommasani, Colin Raffel, Barret Zoph, Sebastian Borgeaud, Dani Yogatama, Maarten Bosma, Denny Zhou, Donald Metzler, Ed Huai hsin Chi, Tatsunori Hashimoto, Oriol Vinyals, Percy Liang, Jeff Dean, and William Fedus. 2022a. Emergent abilities of large language models. Trans. Mach. Learn. Res., 2022.
  • Wei et al. (2022b) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed H Chi, Quoc V Le, Denny Zhou, et al. 2022b. Chain-of-thought prompting elicits reasoning in large language models.
  • Wu et al. (2023) Yiran Wu, Feiran Jia, Shaokun Zhang, Qingyun Wu, Hangyu Li, Erkang Zhu, Yue Wang, Yin Tat Lee, Richard Peng, and Chi Wang. 2023. An empirical study on challenging math problem solving with gpt-4. ArXiv.
  • Xi et al. (2023) Zhiheng Xi, Wenxiang Chen, Xin Guo, Wei He, Yiwen Ding, Boyang Hong, Ming Zhang, Junzhe Wang, Senjie Jin, Enyu Zhou, Rui Zheng, Xiaoran Fan, Xiao Wang, Limao Xiong, Qin Liu, Yuhao Zhou, Weiran Wang, Changhao Jiang, Yicheng Zou, Xiangyang Liu, Zhangyue Yin, Shihan Dou, Rongxiang Weng, Wensen Cheng, Qi Zhang, Wenjuan Qin, Yongyan Zheng, Xipeng Qiu, Xuanjing Huan, and Tao Gui. 2023. The rise and potential of large language model based agents: A survey.
  • Yao et al. (2023) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas L. Griffiths, Yuan Cao, and Karthik Narasimhan. 2023. Tree of thoughts: Deliberate problem solving with large language models. ArXiv, abs/2305.10601.
  • Zelikman et al. (2022) Eric Zelikman, Yuhuai Wu, Jesse Mu, and Noah Goodman. 2022. Star: Bootstrapping reasoning with reasoning. Advances in Neural Information Processing Systems, 35:15476–15488.
  • Zhang et al. (2022a) Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui Chen, Christopher Dewan, Mona Diab, Xian Li, Xi Victoria Lin, Todor Mihaylov, Myle Ott, Sam Shleifer, Kurt Shuster, Daniel Simig, Punit Singh Koura, Anjali Sridhar, Tianlu Wang, and Luke Zettlemoyer. 2022a. Opt: Open pre-trained transformer language models. ArXiv, abs/2205.01068.
  • Zhang et al. (2022b) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alexander J. Smola. 2022b. Automatic chain of thought prompting in large language models. ArXiv, abs/2210.03493.
  • Zhang et al. (2023) Zhuosheng Zhang, Aston Zhang, Mu Li, Hai Zhao, George Karypis, and Alex Smola. 2023. Multimodal chain-of-thought reasoning in language models. arXiv preprint arXiv:2302.00923.
  • Zheng et al. (2023) Chuanyang Zheng, Zhengying Liu, Enze Xie, Zhenguo Li, and Yu Li. 2023. Progressive-hint prompting improves reasoning in large language models. ArXiv, abs/2304.09797.
  • Zhong et al. (2022) Wanjun Zhong, Siyuan Wang, Duyu Tang, Zenan Xu, Daya Guo, Yining Chen, Jiahai Wang, Jian Yin, Ming Zhou, and Nan Duan. 2022. Analytical reasoning of text. In NAACL-HLT.
  • Zhou et al. (2022a) Denny Zhou, Nathanael Scharli, Le Hou, Jason Wei, Nathan Scales, Xuezhi Wang, Dale Schuurmans, Olivier Bousquet, Quoc Le, and Ed Huai hsin Chi. 2022a. Least-to-most prompting enables complex reasoning in large language models. ArXiv, abs/2205.10625.
  • Zhou et al. (2022b) Yongchao Zhou, Andrei Ioan Muresanu, Ziwen Han, Keiran Paster, Silviu Pitis, Harris Chan, and Jimmy Ba. 2022b. Large language models are human-level prompt engineers. ArXiv, abs/2211.01910.
Appendix A Prompt Details
Label
Prompt
Accuracy
C0
Think it step by step.
80.0
C1
Solve the question step by step
84.0
P1
Decompose the given question into smaller segments, elucidating each segment as you rephrase it.
87.0
P2
Break down the following question into concise phrases and elaborate on each phrase while rewriting.
86.0
P3
Rewrite the following question by decomposing it into shorter clauses and providing explanations for each clause.
85.0
P4
Restructure the subsequent question by dissecting it into more concise clauses and enhancing clarity through explanatory rephrasing.
80.0
P5
Break down the problem into independent, concise, and complete phrases, aligning the meaning of each phrase with the original text. Focus on expressing only one concept, action, or condition in each phrase. Then, provide detailed explanations for each phrase, analyzing the implicit messages, defining terms, and using precise professional vocabulary to accurately convey the meaning, aiming to match the potential intention of the target problem.
62.0
Table 11: Results of prompt selection. While P1 was selected and employed in our experiments to demonstrate the effectiveness of the concept of elaboration, there could be better prompts.
Prompt Design and Selection.
We randomly sampled 200 points from GSM8k dataset for prompt selections. The alternatives are generated by communicating with ChatGPT. We test the original zero-shot instruction and the standard CoT used in our experiments. Then, we test four brief prompts and one particularly detailed prompt. The results are list in Table 11. An interesting phenomenon is that the overly detailed prompt (P5) resulted in a significant decrease.
The other prompts used in our experiments.
[…]
All prompts are placed before the given problem. Then we extracted the results from outputs based on regular expressions. However, the outputs may still be diverse and not compliant with the formatting. Therefore, we manually compared and extracted all the non-numeric and non-option items.
Appendix B Example Input and Output Pairs
Q: There were 7 roses in the vase. Jessica cut some more roses from her flower garden which had a total of 59 roses. There are now 20 roses in the vase. How many roses did she cut?
Standard CoT (Solve the question step by step.):
A: To solve the problem, we need to find out how many roses Jessica cut from her garden.
We know that there were originally 7 roses in the vase, and now there are 20. So Jessica must have added 20 - 7 = 13 roses to the vase.
We also know that Jessica cut these roses from her flower garden, which had a total of 59 roses. So the number of roses she cut is: 59 - 13 = 46
Therefore, Jessica cut 46 roses from her garden. (wrong)
PEP + CoT (Decompose the given question into smaller segments, elucidating each segment as you rephrase it. Then, solve the question step by step.):
Segment 1: How many roses were in the vase initially?
- There were 7 roses in the vase.
Segment 2: How many roses did Jessica cut from her flower garden?
- Jessica cut some more roses from her flower garden.
Segment 3: How many roses were in Jessica’s flower garden initially?
- Jessica’s flower garden had a total of 59 roses.
Segment 4: How many roses are in the vase now?
There are now 20 roses in the vase.
Solution:
To find out how many roses Jessica cut from her flower garden, we need to subtract the initial number of roses in the vase from the current number of roses in the vase.
Initial number of roses in the vase = 7
Current number of roses in the vase = 20
Number of roses Jessica added to the vase = Current number of roses in the vase - Initial number of roses in the vase= 20 - 7 = 13
Therefore, Jessica cut 13 roses from her flower garden.
Table 12: The complete example of Fig. 1 from SVAMP, generated by turbo with greedy decoding. The sentence (cyan) is complex and confusing as the misleading condition is mixed with the aiming condition. The standard CoT does not correctly parse the problem (red), while PEP  (blue) benefits from the decomposition and elucidation to clarify the problem.
Question 1:
It is approximately 1955 kilometers from San Diego, California to New York City, New York. If Bernice drove 325 kilometers for 4 days, how many kilometers will she still need to drive?
CoT
To find the total distance Bernice drove, we can multiply the distance she drove each day (325 kilometers) by the number of days she drove (4 days): 325 km/day x 4 days = 1300 kilometers. Therefore, Bernice drove 1300 kilometers. To find how many kilometers she still needs to drive, we can subtract the distance she drove from the total distance between San Diego and New York City: 1955 km - 1300 km = 655 kilometers. Therefore, Bernice still needs to drive 655 kilometers. (Correct)
Error: The decomposition broke the continuous context into segments, changing the related implicit implicatures.
PEP + CoT:
Segment 1: What is the distance between San Diego, California and New York City, New York? The distance between San Diego, California and New York City, New York is approximately 1955 kilometers. Segment 2: How many kilometers did Bernice drive? Bernice drove 325 kilometers. Segment 3: How many days did Bernice drive for? Bernice drove for 4 days. Segment 4: How many kilometers will Bernice still need to drive? To find out how many kilometers Bernice will still need to drive, we need to subtract the distance she has already driven from the total distance between San Diego and New York. 1955 km - 325 km = 1630 km. Therefore, Bernice will still need to drive 1630 kilometers.
Question 2:
Sandy wants to lose as much weight as Joey does but needs 4 weeks to lose the same amount of weight that Joey loses in a single week. If Joey loses 8 pounds in 4 weeks, how many weeks will it take Sandy to lose the same amount of weight?
CoT
First, we need to find out how much weight Joey loses in a single week. We can do this by dividing the total weight he loses in 4 weeks by 4: 8 pounds ÷ 4 weeks = 2 pounds per week. So Joey loses 2 pounds per week. Next, we need to figure out how much weight Sandy wants to lose. We know she wants to lose the same amount as Joey, which is 8 pounds. Now we can use the information given in the problem to find out how many weeks it will take Sandy to lose 8 pounds. We know that Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week. So we can set up a proportion: 4 weeks to lose 2 pounds = x weeks to lose 8 pounds. To solve for x, we can cross-multiply: 4 weeks * 8 pounds = 2 pounds * x weeks; 32 pounds = 2x; Divide both sides by 2: 16 weeks = x. So it will take Sandy 16 weeks to lose the same amount of weight that Joey loses in 4 weeks. (Correct)
Error: The elucidation emphasized and rephrased the given condition, changing the underlying implicatures.
PEP + CoT:
Segment 1: Sandy wants to lose as much weight as Joey does. Elucidation: Sandy has a weight loss goal that is equivalent to Joey’s weight loss goal. Segment 2: Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week. Elucidation: Sandy’s weight loss rate is slower than Joey’s weight loss rate. It takes Sandy 4 weeks to lose the same amount of weight that Joey loses in just one week. Segment 3: Joey loses 8 pounds in 4 weeks. Elucidation: Joey’s weight loss goal is to lose 8 pounds, and he achieves this goal in 4 weeks. Segment 4: How many weeks will it take Sandy to lose the same amount of weight? Elucidation: The question is asking for the number of weeks it will take Sandy to achieve the same weight loss goal as Joey.
Solution: Since Joey loses 8 pounds in 4 weeks, his weight loss rate is 8/4 = 2 pounds per week. Since Sandy needs 4 weeks to lose the same amount of weight that Joey loses in a single week, her weight loss rate is 1/4 = 0.25 pounds per week. To lose 8 pounds like Joey, Sandy will need to lose at the same rate of 2 pounds per week. Therefore, it will take Sandy 8/2 = 4 weeks to lose the same amount of weight as Joey.
Question 3:
Patty’s Plumbing charges \$40 to visit a house to make a repair, plus \$35 per hour, or part thereof, for labor, plus parts. One job took 2.25 hours and used \$60 in parts. How much did Patty charge?
CoT
1. Calculate the labor cost: The job took 2.25 hours, so Patty charged for 3 hours (since she charges for any part of an hour). The labor cost is \$35 per hour, so the labor cost for 3 hours is 3 x \$35 = \$105. 2. Calculate the total cost: The visit fee is \$40. The labor cost is \$105. The parts cost is \$60. The total cost is \$40 + \$105 + \$60 = \$205. Therefore, Patty charged \$205 for the job. (Correct)
Error: PEP focuses on known information and may constrain the associative thinking of CoT during reasoning.
PEP + CoT:
Segment 1: What is the base fee for Patty’s Plumbing to visit a house for a repair? Patty’s Plumbing charges \$40 to visit a house for a repair. Segment 2: How much does Patty’s Plumbing charge for labor? Patty’s Plumbing charges \$35 per hour, or part thereof, for labor. Segment 3: How much does Patty’s Plumbing charge for parts? The job used \$60 in parts. Segment 4: How long did the job take? The job took 2.25 hours.
Solution: To calculate the total cost of the job, we need to add the base fee, labor cost, and parts cost. Base fee: \$40 Labor cost: \$35 per hour x 2.25 hours = \$78.75 Parts cost: \$60. Total cost: \$40 + \$78.75 + \$60 = \$178.75. Therefore, Patty charged \$178.75 for the job.
Table 13: Error cases from GSM8k. For certain sentences in the problem (cyan), CoT can understand it correctly (blue), while PEP may lead to erroneous reasoning and generate errors (red).
The complete results and comparison of Fig. 1 is shown in Table 12. The given sentence in the problem is complex and confusing, as it mixed the condition "Jessica cut some more roses" which is associated with the final question and the irrelevant sentence "which had a total of 59 roses". CoT is easily misled, while PEP correctly decomposes these conditions into two segments and elucidate the behind meanings, where the downstream reasoning process gets more clear and coherent.
In Table 13 we provide more error cases. As shown, LLMs could also make mistakes during elaboration, misleading the following reasoning. PEP may break down the continuous context, thus changing the implicit implicatures. Besides, PEP may focus too much on the locality of sentences, constraining the associative thinking of CoT.